# ESGenius: Benchmarking LLMs on Environmental, Social, and Governance (ESG) and Sustainability Knowledge

### v1

@@ 4.2.7 Interactive Visualization
Figure 7 presents an interactive heat-map visualization of model performance across 50 LLMs (ranging from 0.5 to 671 billion parameters) and all 1,136 MCQs in ESGenius. The visualization arranges models (rows) by their overall zero-shot accuracy and questions (columns) by difficulty, creating a clear left-to-right progression from predominantly red cells (incorrect/unanswered questions) to green cells (correct answers). Interactive features enhance analysis - hovering over any cell reveals detailed information including the model name and rank, question ID, difficulty score, complete prompt, answer options, ground truth, and the model’s prediction. An example of this interactive tooltip is shown in Figure 8. This dynamic visualization transforms raw accuracy data into an intuitive diagnostic tool for analyzing model performance patterns and systematic weaknesses. The interactive visualization is publicly accessible at https://angel-ntu.github.io/ESGenius/.
@@ 5 Limitations
While ESGenius aspires to provide a holistic evaluation of ESG understanding, several caveats must be acknowledged:
@@
  • Coverage trade-offs——the benchmark covers only seven sources, which may not comprehensively cover all ESG topics and sustainability issues. While these sources represent key frameworks and standards, they exclude many specialized industry guidelines, regional regulations, and emerging sustainability frameworks.
  • Expert dependency—crafting high-fidelity questions and validating answers hinges on scarce domain specialists, which introduces both capacity limits and potential disciplinary bias. The specialized nature of ESG expertise makes scaling the question creation process especially difficult.
@@
  • Format constraints—standardized multiple-choice questions facilitate large-scale scoring but compress the nuanced reasoning and synthesis often required in real-world ESG analysis. This format may not fully capture the complexity of real-world sustainability decision-making.
@@
  • Standard volatility—frameworks such as ISSB guidance and IPCC assessments evolve quickly, obliging continual corpus refreshes and version control to stay current. This rapid evolution makes maintaining benchmark relevance an ongoing challenge.
  • Limited language—the present focus on English, globally dominant standards risks under-representing regional regulations and non-English corpora, potentially missing important local sustainability contexts and practices.
  • Modality gaps—the text-only design excludes visual disclosures (e.g., charts, geospatial maps) and time-series data that are central to many ESG reports, limiting the benchmark’s ability to test comprehension of multi-modalities.
  • Simplified scoring—binary accuracy overlooks partial correctness and fails to credit sophisticated but incomplete reasoning, calling for finer-grained metrics that can better capture nuanced understanding.
@@
  • Copyright limits—licensing restrictions preclude the inclusion of certain proprietary ESG standards, thereby limiting the potential for truly exhaustive coverage and potentially missing key industry-specific frameworks.
6 Conclusion
We presented ESGenius, the first benchmark dedicated to probing LLMs on the breadth and depth of ESG and sustainability knowledge, unifying authoritative standards, corporate questionnaires, and scientific assessments into a multi-dimensional QA dataset. A two-stage evaluation protocol—Zero-Shot followed by RAG—reveals that grounding responses in curated evidence boosts accuracy by roughly 15–30 percentage points, and that 4 to 72 billion parameter models equipped with targeted knowledge can surpass much larger zero-shot models, underscoring the primacy of domain integration over sheer scale. Looking ahead, we aim to incorporate emerging regulations, automate corpus refreshes, extend coverage to non-English sources, and design finer-grained metrics that capture partial credit and nuanced ESG reasoning. Overall, ESGenius is designed to serve as a dedicated benchmark to facilitate the development of sustainability knowledge-aware AI systems in the complex ESG domain.
@@ A.1 Automated Preparation of Candidate MCQs
This appendix details the fully-automated preparation stage that converts raw knowledge sources into a pool of candidate multiple-choice questions (MCQs). (The complete algorithm is detailed in Algorithm 1.) These candidate questions subsequently undergo expert review, editing, and validation before potential inclusion in ESGenius-QA. The goals of the preparation stage are twofold: (i) breadth—to cover as many distinct passages as possible from the source corpus, and (ii) difficulty—to generate questions that force LLMs to reason beyond surface-level facts.
@@
optionally concatenating the next page when the combined length remains under the upper bound. Each chunk is tagged with its page span (e.g. “12–13”) so that provenance is preserved.
@@
For every selected chunk the pipeline sends a carefully engineered prompt (see A) to the Qwen Max’s DashScope endpoint111qwen-max-2025-01-25, temperature 0, deterministic for reproducibility. that requests exactly one “extremely challenging” MCQ. Mandatory instructions enforce (i) deep reasoning across multiple sentences, (ii) near-miss distractors, (iii) answerability from the excerpt alone, and (iv) strict JSON output. A fixed seed (42) controls chunk sampling so that runs are reproducible.
@@
The returned JSON is sanitised, parsed, and validated. Missing or malformed keys trigger rejection. The correct answer text is cross-checked with the declared key, and a fallback correction is applied if necessary. A universal Z option (“Not sure”) is appended to every item to enable abstention analysis. Each validated MCQ is wrapped in a metadata envelope containing {id, source_pdf, page_range, difficulty, generation_timestamp}.
@@
Consistency checking and question de-duplication.
@@
The resulting candidate pool is then passed to the expert review process detailed in Section 3.1. This marks the transition from automated preparation to rigorous human validation and refinement of each question.
@@
  "correct_option_key": "String: The key (’A’, ’B’, ’C’, or ’D’) corresponding to the correct answer."
@@ A.3 ESGenius-QA Example Question Structures
Zero-shot Evaluation Example
@@
Annual mean concentration of PM2.5 (µg/m³) 7.4 (2022) •D
@@
CO2 emissions from fossil-fuel combustion and cement production (t CO2/capita) 1.2 (2022) •A
GHG emissions embodied in imports (t CO2/capita) NA NA ••
CO2 emissions embodied in fossil-fuel exports (kg/capita) 0.0 (2022) ••
@@ Appendix B Prompt Template
For multiple-choice questions (4 options + 1 Not sure with 1 answer), we use the following prompt templates:
@@ Appendix C Main Experimental Results Table
Alibaba Qwen Qwen2.5-Max 22B†(MoE 325B) XL No Yes 0.6444 - -
@@
Table 2: Comprehensive ESGenius results showing LLM performance under Zero-shot and RAG settings. S.G. denotes Size Group (S: Small ($\leq$1B), M: Medium (1–7B), L: Large (7–30B), XL: Extra Large (30–100B), XXL: Extra Extra Large (>>100B)); I-T: Instruction Tuning; Rea: Reasoning Focus. Improvement (%): $\frac{\text{RAG}-\text{Zero-shot}}{\text{Zero-shot}}\times 100$. “–” indicates scores currently unavailable due to technical constraints. † indicates industry size estimates.
@@ Appendix D ESGenius-Corpus
The corpus integrates authoritative frameworks, corporate questionnaires, and scientific assessments that comprehensively cover environmental, social, and governance (ESG) considerations. A detailed quantitative overview of the collection is presented across multiple tables: Table 3 summarizes IPCC climate-science assessments, Table 4 covers GRI Standards, Tables 56 detail SASB’s industry-specific disclosures, Tables 78 outline IFRS/ISSB standards and guidance, Table 9 catalogs TCFD materials, Table 10 documents CDP questionnaires, and Table 11 captures UN Sustainable Development Goals content. The distribution of pages and questions across these sources is visualized in Figures 5 and 6, respectively. From this comprehensive dataset, several key insights emerge:
@@
    Sustainable development anchor. The UN’s Sustainable Development Goals (SDGs) underpin cross-sector and cross-country sustainability objectives. Although comparatively compact in page count, these seminal UN publications situate corporate ESG strategies within the global 2030 Agenda, ensuring broader alignment with international development priorities.
@@
    Imbalanced density highlights practical challenges. Marked disparities between the distribution of documents and the distribution of pages underscore the varied scope of ESG sources: some (e.g. IPCC) are exhaustive scientific compendiums, while others (e.g. GRI, SASB, IFRS) comprise slimmer but more numerous reference standards. For researchers and practitioners alike, tasks ranging from large-scale summarization to specialized technical queries must navigate this imbalance of item count versus depth.
@@
Table 3: Comprehensive metadata for the seven IPCC reports curated in the ESGenius. This collection represents the complete Sixth Assessment Report (AR6) cycle and key Special Reports from 2018-2023. The AR6 materials include the 2023 Synthesis Report and three Working Group contributions covering physical science (WG1), impacts & adaptation (WG2), and mitigation (WG3). Three thematic Special Reports address land use (SRCCL), oceans & ice (SROCC), and 1.5°C warming pathways (SR15). Totaling 10,582 pages and 117 evaluation questions, these authoritative climate science assessments form a crucial knowledge foundation for ESG analysis. Each report is with document sizes ranging from 4.9MB to 378MB. The collection provides comprehensive coverage of climate science, impacts, and policy responses that inform modern ESG frameworks like TCFD, CSRD, and ISSB standards. All documents are sourced directly from IPCC (https://www.ipcc.ch) and represent peer-reviewed, UN-mandated scientific assessments.
@@
Table 6: Comprehensive metadata for all 77 industry-specific SASB Sustainability Accounting Standards issued in 2023 and now stewarded by the IFRS Foundation’s ISSB. This corpus, distilled for our ESGenius system, comprises 77 documents totalling 30.431 MB, 1,864 pages, and 236 multiple-choice questions. The standards span the full economy—from Apparel to Transportation—and cover financially material sustainability topics including greenhouse gas emissions, water management, data security, and workforce diversity. These standards provide investors with decision-relevant ESG information for valuation, risk assessment, and stewardship. The complete collection is freely accessible at https://www.sasb.org/.
@@
Table 8: Comprehensive metadata for the entire IFRS / ISSB sustainability-disclosure corpus (2023–2025): it comprises the universal core standards IFRS S1 (general sustainability-related financial disclosure) and IFRS S2 (climate-related disclosure), their bases for conclusions, accompanying guidance, the 2024 climate-progress report, the 2024 Sustainability Disclosure Taxonomy, the 2025 greenhouse-gas exposure draft (with basis), a comparison of IFRS S2 with the TCFD recommendations, and the 68-volume IFRS S2 Industry-based Guidance that maps SASB’s sector-specific materiality into the new framework. Together the 77 PDFs, total 20.81 MB, span 1 866 pages—well over two thousand when including ancillary matter—and yield 238 benchmark MCQs for ESGenius. IFRS S1 sets the universal disclosure baseline, while IFRS S2 details climate-specific metrics, mirroring the TCFD architecture and enriched with SASB’s sectoral depth; the accompanying materials establish a globally consistent, investor-focused baseline that links decision-relevant sustainability information—such as greenhouse-gas emissions, transition plans, climate resilience, data security, and workforce diversity—directly to financial statements. All documents are freely available at https://www.ifrs.org/.
@@
Figure 6: Relative page‐count distribution of the ESGenius-Corpus across its 7 source families. The Intergovernmental Panel on Climate Change (IPCC) alone accounts for a majority of the material—54 %, or 10 582 pages—reflecting the encyclopaedic scope of its assessment reports. Standards and disclosure frameworks contribute smaller but still substantive shares: CDP (10.1 %, 1 987 pp.), IFRS/ISSB (9.5 %, 1 866 pp.), SASB (9.5 %, 1 864 pp.), and GRI (9.3 %, 1 826 pp.). Guidance‐oriented sources such as TCFD (3.7 %, 728 pp.) and the UN Sustainable Development Goals corpus (3.8 %, 735 pp.) round out the collection.
Figure 7: Example zero-shot performance heatmap showing model accuracy patterns across different ESG question types and topics within ESGenius (based on initial data). Darker colors indicate higher accuracy. Interactive version url will be provided in the camera-ready version.

### v2

@@ 4.2.8 Interactive Visualization
Figure 9 presents an interactive heatmap visualization of model performance across 50 LLMs (ranging from 0.5 to 671 billion parameters) and all 1,136 MCQs in ESGenius. The visualization arranges models (rows) by their overall zero-shot accuracy and questions (columns) by difficulty, creating a clear left-to-right progression from predominantly red cells (incorrect/unanswered questions) to green cells (correct answers). Interactive features enhance analysis - hovering over any cell reveals detailed information including the model name and rank, question ID, difficulty score, complete prompt, answer options, ground truth, and the model’s prediction. An example of this interactive tooltip is shown in Figure 10. This dynamic visualization transforms raw accuracy data into an intuitive diagnostic tool for analyzing model performance patterns and systematic weaknesses. The interactive visualization is publicly accessible at https://angel-ntu.github.io/ESGenius.
5 Conclusion
We presented ESGenius, the first comprehensive QA benchmark dedicated to probing LLMs on the breadth and depth of ESG and sustainability knowledge. It unifies two tightly integrated components: ESGenius-QA, a large-scale, expert-validated MCQs dataset, and the ESGenius-Corpus, a curated and extensible collection of authoritative ESG sources that anchors every question with verifiable evidence. Our two-stage evaluation protocol—Zero-Shot followed by RAG—shows that grounding responses in curated evidence boosts accuracy by 15–30 percentage points, with smaller models enhanced by RAG often surpassing much larger zero-shot models. This demonstrates that domain integration and transparent grounding are more decisive than sheer scale alone. By open-sourcing data, code, and a live and interactive leaderboard, ESGenius establishes a reproducible and continually updated yardstick for sustainability-aware AI. Our subsequent work will expand this mission to multi-modal ESG tasks Zhang et al. (2025), alongside expert-in-the-loop auditing, paving the way toward trustworthy ESG decision support.
@@ 6 Limitations
While ESGenius aspires to provide a holistic evaluation of ESG understanding, several caveats must be acknowledged.
Coverage trade-offs.
@@
The benchmark currently includes only seven sources, which, while representing key frameworks and standards, cannot comprehensively encompass all ESG topics, industry-specific guidelines, regional regulations, or emerging sustainability frameworks.
Expert dependency and scalability.
The creation of high-fidelity questions and the validation of answers are heavily dependent on scarce domain experts, making the process capacity-limited and prone to disciplinary bias. In future iterations, we plan to explore scalable strategies such as expert-in-the-loop active sampling or agreement-based adjudication to sustain long-term benchmark growth.
Format constraints.
@@
The reliance on standardized MCQs, although enabling large-scale scoring, compresses the nuanced reasoning and synthesis required in real-world ESG analysis and may therefore fall short of fully reflecting the complexity of sustainability decision-making.
Standard volatility.
@@
Frameworks such as ISSB guidance and IPCC assessments evolve rapidly, necessitating continual updates, corpus refreshes, and careful version control to maintain relevance.
Language limitations.
The current focus on English-only corpora risks under-representing non-English ESG documents, local regulations, and region-specific sustainability practices. Expanding coverage to multilingual corpora is a key priority for improving global relevance.
Exclusion of visual elements.
ESG reports often include key visual disclosures—charts, graphs, maps, and time-series plots. Our text-only design excludes multimodal reasoning. While text remains central, future versions will integrate visuals via OCR pipelines and vision–language models to better capture ESG reporting.
Metrics and evaluation scope.
Reliance on binary accuracy misses partial correctness and deeper ESG reasoning. We plan to add metrics such as explanation consistency, factual groundedness, and chain-of-thought plausibility. The MCQ format, while limited, offers scalability and comparability, as seen in MMLU and MMMU. We are exploring richer formats—open-ended answers, reasoning traces, and retrieval tasks—for better alignment with real-world ESG analysis.
Bias in LLM-generated questions.
LLMs are widely used in benchmarks for question generation and evaluation. Our pipeline adds expert verification to ensure factuality and relevance, but both models and humans can introduce bias. While we emphasize consistency and oversight, residual biases remain. We welcome validated alternatives to further reduce bias in future versions.
Limitations of the RAG implementation.
We use a simple RAG setup for reproducibility, deployment with smaller LLMs, and clear attribution of performance gains to the dataset and task design. Despite its simplicity, it already yields strong improvements. Future versions will integrate more advanced retrieval to enhance robustness and real-world alignment.
Human baseline performance.
Without human benchmarks, model scores lack context. We are developing a crowdsourcing interface for experts to provide evaluations, establishing baselines that will ground interpretation of results and guide future benchmarking.
Copyright limits.
@@
Licensing restrictions constrain the inclusion of certain proprietary ESG standards, limiting exhaustive coverage and potentially omitting significant industry-specific frameworks.
7 Ethical Considerations
Purpose and scope.
ESGenius is released to advance trustworthy, evidence–grounded research on LLMs Environmental, Social, and Governance (ESG) question answering.
Sources, copyright, and licensing.
The ESGenius-Corpus aggregates authoritative but heterogeneous materials (e.g., GRI, SASB/ISSB, IPCC, TCFD, CDP, SDGs in §3.2). Many of these works are copyrighted and/or distributed under specific use terms. We therefore: (i) include only publicly available documents or clearly reference official distribution pages; (ii) link every question to minimal excerpts sufficient for verification, avoiding wholesale redistribution; and (iii) ship documentation that records provenance and access routes for each source. Users of ESGenius must review and comply with the licenses and terms of the original sources; the benchmark is provided strictly for academic research and reproducibility.
LLM assistance and data leakage.
LLMs were used only to propose candidate MCQs; six domain experts double–validated all retained questions and three senior practitioners consulted the process (§3.1). To reduce circularity, we (i) enforced strict evidence requirements per question, (ii) rejected or revised questions that were not unambiguously supported, and (iii) evaluated models—including the assisting model—via API in an isolated session, with no access to construction traces. Each question is mapped to source passages (§3.1), and a “Not sure (Z)” option was added to discourage hallucination and reward calibrated abstention (§3.1). Beyond content generation, LLMs were occasionally employed as supportive tools for surface-level polishing, such as refining phrasing in writing or improving code readability and consistency. These uses were strictly non-substantive and did not influence the design, validation, or evaluation of ESGenius itself.
Annotation practice and labor.
All validators were experienced in sustainability or NLP and gave informed consent to participate. The workflow emphasized factual verification, clarity, and distractor plausibility, with systematic rejection of ambiguous questions (§3.1). We credit contributors as co–authors, and we encourage future adopters to follow fair labor practices when extending the dataset.
Fairness, representation, and epistemic balance.
ESGenius currently emphasizes English–language, investor–oriented frameworks and global standards (e.g., IFRS S1/S2, SASB) over national, local, Indigenous, or multilingual sources (§6). This may reproduce an institutional or Global–North lens and under–represent region–specific norms or community priorities. We flag this as an ethical limitation and design the corpus to be modular and extensible3.2). Users should avoid over–generalizing conclusions beyond the covered sources and geographies.
Dual–use risks and misuse.
Benchmarking can be gamed or misapplied. Examples include (i) training to the test (overfitting to known questions) and (ii) using ESG QA capability claims to enable persuasive but ungrounded communication or greenwashing. To mitigate these risks we: (a) link every question to authoritative evidence and encourage evaluation settings that require citations (RAG, §4.1.2); (b) provide versioning and clear changelogs; (c) retain the “Z” option to penalize confident guessing; and (d) discourage using ESGenius scores as sole evidence of real–world compliance or as a checklist for circumventing disclosure obligations. We further recommend reporting retrieval quality and evidence coverage alongside accuracy.
Leaderboard and claims.
Model rankings can invite outsized or decontextualized claims. We caution that (i) ESGenius is one lens among many (§6); (ii) zero–shot accuracy reflects parametric knowledge, whereas RAG performance reflects retrieval and grounding design (§4.2); and (iii) differences of a few percentage points may not be practically meaningful without error and subtopic analyses (§4.2.7). Public results should disclose prompts, decoding settings, retrieval parameters, and version identifiers to support comparability and reproducibility.
Safety of model outputs.
ESG advice or interpretations can carry legal, financial, or societal consequences. ESGenius is not a substitute for professional guidance. We encourage deployers to (i) present grounded citations by default; (ii) surface uncertainty via calibrated abstention; and (iii) keep a expert-in-the-loop for consequential uses (e.g., risk management, regulatory filings). Our findings (e.g., stronger gains with RAG) underscore that evidence-linked responses are ethically preferable to unsupported generations.
Environmental impact.
Benchmarking LLMs consumes energy and incurs carbon costs. Our experiments used a single DGX A100 node (4 $\times$ 80GB) to limit overhead. We encourage (i) reporting hardware and token budgets, (ii) batching and caching retrieval, and (iii) preferring smaller models with strong RAG performance when feasible (§4.2.1).
Responsible release and governance.
We release ESGenius with (i) documented data lineage and versioning; (ii) explicit usage guidelines discouraging compliance automation or deceptive communication; and (iii) maintenance plans for updates as standards evolve (§3.2, §5). Community feedback on coverage gaps, bias, and error reports is welcome, and validated contributions will be incorporated through transparent review.
@@ A.1 Automated Preparation of Candidate MCQs
This appendix details the fully-automated preparation stage that converts raw knowledge sources into a pool of candidate MCQs. (The complete algorithm is detailed in Algorithm 1.) These candidate questions subsequently undergo expert review, editing, and validation before potential inclusion in ESGenius-QA. The goals of the preparation stage are twofold: (i) breadth—to cover as many distinct passages as possible from the source corpus, and (ii) difficulty—to generate questions that force LLMs to reason beyond surface-level facts.
@@
optionally concatenating the next page when the combined length remains under the upper bound. Each chunk is tagged with its page span (e.g., “12–13”) so that provenance is preserved.
@@
For every selected chunk the pipeline sends a carefully engineered prompt (§A) to the Qwen Max’s DashScope endpoint111qwen-max-2025-01-25, temperature 0, deterministic for reproducibility. that requests exactly one “extremely challenging” MCQ. Mandatory instructions enforce (i) deep reasoning across multiple sentences, (ii) near-miss distractors, (iii) answerability from the excerpt alone, and (iv) strict JSON output. A fixed seed (42) controls chunk sampling so that runs are reproducible.
@@
The returned JSON is sanitised, parsed, and validated. Missing or malformed keys trigger rejection. The correct answer text is cross-checked with the declared key, and a fallback correction is applied if necessary. A universal Z option (“Not sure”) is appended to every question to enable abstention analysis. Each validated MCQ is wrapped in a metadata envelope containing {id, source_pdf, page_range, difficulty, generation_timestamp}.
@@
Consistency checking and question deduplication.
@@
The resulting candidate pool is then passed to the expert review process detailed in §3.1. This marks the transition from automated preparation to rigorous human validation and refinement of each question.
@@
  "correct_option_key": "String: The key (’A’, ’B’, or ’C’, or ’D’) corresponding to the correct answer."
@@ A.3 ESGenius-QA Example Question Structures
A.3.1 Zero-Shot Evaluation Example
@@ A.3.2 RAG Evaluation Example
Annual mean concentration of $PM_{2.5}$ (µg/m³) 7.4 (2022) •D
@@
$CO_{2}$
emissions from fossil-fuel combustion and cement production (t $CO_{2}$/capita) 1.2 (2022) •A
GHG emissions embodied in imports (t $CO_{2}$/capita) NA NA ••
$CO_{2}$ emissions embodied in fossil-fuel exports (kg/capita) 0.0 (2022) ••
@@ Appendix B Prompt Template
For MCQs (4 options + 1 Not sure with 1 answer), we use the following prompt templates:
@@ Appendix C Main Experimental Results Table
Alibaba Qwen Qwen2.5-Max (qwen-max-2025-01-25) A22B†(MoE 325B) XL No Yes 0.6444 - -
@@
Table 2: Comprehensive ESGenius results showing LLM performance under Zero-Shot and RAG settings. S.G. denotes Size Group (S: Small ($\leq$1B), M: Medium (1–7B), L: Large (7–30B), XL: Extra Large (30–100B), XXL: Extra Extra Large (>>100B)); I-T: Instruction Tuning; Rea: Reasoning Focus. Improvement (%): $\frac{\text{RAG}-\text{Zero-Shot}}{\text{Zero-Shot}}\times 100$. “–” indicates scores currently unavailable due to technical constraints. † indicates industry size estimates. A ranking bar chart of zero-shot performance is shown in Figure 3.
Figure 3: Zero-Shot model ranking on ESGenius (1136 questions). The chart shows model accuracies across 50 LLMs, with o3 achieving the highest score of 72.54%.
@@ Appendix D ESGenius-Corpus
The corpus integrates authoritative frameworks, corporate questionnaires, and scientific assessments that comprehensively cover environmental, social, and governance (ESG) considerations. A detailed quantitative overview of the collection is presented across multiple tables: Table 3 summarizes IPCC climate-science assessments, Table 4 covers GRI Standards, Tables 56 detail SASB’s industry-specific disclosures, Tables 78 outline IFRS/ISSB standards and guidance, Table 9 catalogs TCFD materials, Table 10 documents CDP questionnaires, and Table 11 captures UN Sustainable Development Goals (SDGs) content. The distribution of pages and questions across these sources is visualized in Figures 7 and 8, respectively. From this comprehensive dataset, several key insights emerge:
@@
    Sustainable development anchor. The UN’s SDGs underpin cross-sector and cross-country sustainability objectives. Although comparatively compact in page count, these seminal UN publications situate corporate ESG strategies within the global 2030 Agenda, ensuring broader alignment with international development priorities.
@@
    Imbalanced density highlights practical challenges. Marked disparities between the distribution of documents and the distribution of pages underscore the varied scope of ESG sources: some (e.g., IPCC) are exhaustive scientific compendiums, while others (e.g., GRI, SASB, IFRS) comprise slimmer but more numerous reference standards. For researchers and practitioners alike, tasks ranging from large-scale summarization to specialized technical queries must navigate this imbalance of question count versus depth.
@@
Table 3: Comprehensive metadata for the seven IPCC reports curated in the ESGenius. This collection represents the complete Sixth Assessment Report (AR6) cycle and key Special Reports from 2018-2023. The AR6 materials include the 2023 Synthesis Report and three Working Group contributions covering physical science (WG1), impacts & adaptation (WG2), and mitigation (WG3). Three thematic Special Reports address land use (SRCCL), oceans & ice (SROCC), and 1.5°C warming pathways (SR15). Totaling 10,582 pages and 117 evaluation questions, these authoritative climate science assessments form a crucial knowledge foundation for ESG analysis. Document sizes range from 4.9MB to 378MB. The collection provides comprehensive coverage of climate science, impacts, and policy responses that inform modern ESG frameworks like TCFD, CSRD, and ISSB standards. All documents are sourced directly from IPCC (https://www.ipcc.ch) and represent peer-reviewed, UN-mandated scientific assessments.
@@
Table 6: Comprehensive metadata for all 77 industry-specific SASB Sustainability Accounting Standards issued in 2023 and now stewarded by the IFRS Foundation’s ISSB. This corpus, distilled for our ESGenius system, comprises 77 documents totalling 30.431 MB, 1,864 pages, and 236 MCQs. The standards span the full economy—from Apparel to Transportation—and cover financially material sustainability topics including greenhouse gas emissions, water management, data security, and workforce diversity. These standards provide investors with decision-relevant ESG information for valuation, risk assessment, and stewardship. The complete collection is freely accessible at https://www.sasb.org/.
@@
Table 8: Comprehensive metadata for the entire IFRS / ISSB sustainability-disclosure corpus (2023–2025): it comprises the universal core standards IFRS S1 (general sustainability-related financial disclosure) and IFRS S2 (climate-related disclosure), their bases for conclusions, accompanying guidance, the 2024 climate-progress report, the 2024 Sustainability Disclosure Taxonomy, the 2025 greenhouse-gas exposure draft (with basis), a comparison of IFRS S2 with the TCFD recommendations, and the 68-volume IFRS S2 Industry-based Guidance that maps SASB’s sector-specific materiality into the new framework. Together the 77 PDFs, total 20.81 MB, span 1,866 pages—well over two thousand when including ancillary matter—and yield 238 benchmark MCQs for ESGenius. IFRS S1 sets the universal disclosure baseline, while IFRS S2 details climate-specific metrics, mirroring the TCFD architecture and enriched with SASB’s sectoral depth; the accompanying materials establish a globally consistent, investor-focused baseline that links decision-relevant sustainability information—such as greenhouse-gas emissions, transition plans, climate resilience, data security, and workforce diversity—directly to financial statements. All documents are freely available at https://www.ifrs.org/.
@@
Figure 8: Relative page‐count distribution of the ESGenius-Corpus across its 7 source families. The Intergovernmental Panel on Climate Change (IPCC) alone accounts for a majority of the material—54%, or 10,582 pages—reflecting the encyclopaedic scope of its assessment reports. Standards and disclosure frameworks contribute smaller but still substantive shares: CDP (10.1%, 1,987 pp.), IFRS/ISSB (9.5%, 1,866 pp.), SASB (9.5%, 1,864 pp.), and GRI (9.3%, 1,826 pp.). Guidance‐oriented sources such as TCFD (3.7%, 728 pp.) and the UN SDGs corpus (3.8%, 735 pp.) round out the collection.
Figure 9: Example zero-shot performance heatmap showing model accuracy patterns across different ESG question types and topics within ESGenius (based on initial data). Darker colors indicate higher accuracy.